Grade 12Algebra

Determine if the following ratios form a proportion. Also, write the middle terms and extreme terms where the ratios form a proportion.
39 liters: 65 liters and 6 bottles: 10 bottles

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

39 litres : 65 litres and 6 bottles : 10 bottles
39 / 65 = 3 /5
6 / 10 = 3 / 5
Yes, these are in a proportion
Middle terms are 65 litres, 6 bottles and Extreme terms are 39 litres, 10 bottles

Thanks